OpenStudy (anonymous):

someone help me find L!!! S=2lw+2lh+2hw

OpenStudy (radar):

S=l(2w+2h) + 2hw l(2w+2h)=S-2hw l=(S-2hw)/(2w+2h)

OpenStudy (anonymous):

thank you soooo much!! i was so stuck! youre a life saver!